# Benchmarking the Performance of Large Language Models on the Cerebras Wafer Scale Engine

### v1

@@ Abstract
In this paper, we evaluate the performance of LLMs on the Cerebras Wafer Scale Engine (WSE). Cerebras WSE is a high performance computing system with 2.6 trillion transistors, 850,000 cores and 40 GB on-chip memory. Cerebras WSE’s Sparse Linear Algebra Compute (SLAC) cores eliminates multiply-by-zeros operations and its 40 GB of on-chip memory is uniformly distributed among SLAC cores, enabling fast local access to model parameters. Moreover, Cerebras software configures routing between cores at runtime, optimizing communication overhead among cores. As LLMs are becoming more commonly used, new hardware architectures are needed to accelerate LLMs training and inference. We benchmark the effectiveness of this hardware architecture at accelerating LLMs training and inference. Additionally, we analyze if Cerebras WSE can scale the memory-wall associated with traditionally memory-bound compute tasks using its 20 PB/s high bandwidth memory. Furthermore, we examine the performance scalability of Cerebras WSE through a roofline model. By plotting performance metrics against computational intensity, we aim to assess their effectiveness at handling high compute-intensive LLMs training and inference tasks.
@@ II-A Large Language Models (LLMs)
Large language models have gained growing interest over the past few years. Especially with the release of ChatGPT [19], users are impressed with the capabilities of LLMs and can leverage them to assist with daily tasks that were previously unimaginable. Because of its remarkable language understanding abilities, more than one million users signed up to use ChatGPT within one week of its release [20]. Moreover, LLMs have created a profound influence in many industries and fields. For example, LLMs ability to generate and debug code can accelerate the software development process and enhance programmers’ productivity. LLMs can also assist with the creative writing process by offering new ideas and providing feedback to writings [20].
@@
$\bm{A}=\bm{softmax(\frac{{Q}{K}^{T}}{\sqrt[]{{d}_{k}}})}$ (4)
$\bm{{W}^{Q}}\in\mathbb{R}^{d\times{d}_{k}}$, $\bm{{W}^{K}}\in\mathbb{R}^{d\times{d}_{k}}$, $\bm{{W}^{V}}\in\mathbb{R}^{d\times{d}_{V}}$ are the weight matrices for query, key, and value, respectively. $\bm{Q}\in\mathbb{R}^{n\times{d}_{k}}$, $\bm{K}\in\mathbb{R}^{n\times{d}_{k}}$, $\bm{V}\in\mathbb{R}^{n\times{d}_{v}}$ are the query, key, and value matrices, respectively. $\bm{X}\in\mathbb{R}^{n\times d}$ is the input sentence embedding matrix, where $n$ is the input sequence length, $d$ is the embedding length of each word in the input sentence. The SA output $A$ will then be projected to the same shape as input $X$ through a feed-forward network. This sequences of operations define the essential elements of a transformer block and is also called a head. In transformer-based LLMs, each transformer block can contain multiple heads, each head captures different relationships among the input sequence, enabling the model to have a richer context and understanding of the input sequence. This mechanism is called Multi-Head Self Attention (MHSA). The fact that transformer block’s input and output dimensions are the same enables multiple transformer blocks to stack, allowing earlier transformer blocks to provide more context for later blocks.
Some mechanisms where transformer models are trained next word prediction [21] and masked language modeling [22].
@@ II-B GPT
GPT is a transformer decoder-only architecture where each input only computes SA among previous input tokens. Namely, when computing output SA value ${{a}_{i}}$, SA only computes among ${{q}_{i}}$ with ${{k}_{j}}$, ${{v}_{j}}$ where $j\leq i$. This ensures that the model only attends to past and current tokens, not future ones. At each time stamp, GPT computes a probability distribution over the entire vocabulary to predict the probability of next word, enabling it to generate the next word. This process is also called auto-regressive text generation. Usually, auto-regressive text generation until a $|<\text{endoftext}>|$ token is generated or a predefined maximum output sequence length is reached.
Figure 1: GPT transformer decoder architecture. Input embeddings are first added with positional embeddings and then fed into GPT transformer block, where MHSA occurs.
@@ II-C BERT
BERT[12] is a transformer encoder-only architecture where each input computes SA among both previous and future input tokens, called bidirectional SA. Namely, when computing output SA value ${{a}_{i}}$, SA computes among ${{q}_{i}}$, with ${{k}_{j}}$, ${{v}_{j}}$ where $j\leq n$, where $n$ is the input sequence length. Because SA is computed among the entire input sequence. BERT models are able to have a more comprehensive and coherent understanding of the input sentence. Normally, a [CLS] token is added at the beginning of the input sequence. At the output layer, the final hidden state of the [CLS] token is used for the classification of the entire sentence. This ensures that the final classification of the sentence is not biased toward any specific word in the input sequence. BERT models are typically pre-trained on tasks like Masked Language Modeling (MLM) [12], where a random subset of the input sequence is masked and the model predicts the masked words. BERT models are also pre-trained via Next Sentence Prediction, where the model predicts whether 2 given sentences should follow each other. After pre-training, BERT models are fine-tuned for task-specific applications, such as sentiment analysis of input sequence [23].
Figure 2: BERT transformer encoder architecture. [CLS] token embedding is appended to the front of the input sequence. The BERT architecture is similar to the GPT model architecture shown in Figure 1, except that MHSA occurs between the current token and tokens to the left and right. In BERT, the final output hidden state of [CLS] token is used to predict the probability distribution for classification of the input sequence
@@ II-D Cerebras WSE
All training and inference experiments are conducted on the Cerebras Wafer Scale Engine (WSE) and Cerebras CS-2 system. Cererbras WSE is a powerful AI accelerator purposely built to train LLMs [24]. As of May 2023, Cerebras WSE is the largest chip ever built, at 46000 $mm^{2}$ and containing 2.6 trillion transistors and 850,000 cores. Cerebras WSE’s large die size enables greater compute power and enables more computations to occur on-chip [16]. This reduces off-chip communications and enables computation resources to be fully utilized. Cerebras’ chip is built with Taiwan Semiconductor Manufacturing Company (TSMC) 7 $\mu m$ process and runs at 1.1 GHz frequency [15]. Cerebras WSE’s physical design contributes 50% silicon area to static random access memory (SRAM) and 50% to computation logic in each core. Each core features a local 48KB SRAM that enables the core to have fast access to local memory. In each core, memory is organized into eight 6kB banks, each 32-bit wide. On top of the local 48kB memory, each core also has a 256-byte, 64-bit wide, software-configured cache, enabling even faster data access for most frequently used data [15]. This uniformed distributed on-chip memory architecture enables low memory access latency for each core and very high aggregated memory throughput.
@@
Cerebras WSE’s fine-grained data flow scheduling, enables cores to only perform computations on nonzero data, saving dynamic power of cores [15]. The combination of this and the high memory bandwidth enables efficient computations on data with unstructured sparsity [25], which is well suited for neural network computations as model weights and input data can possess arbitrary levels of sparsity. Moreover, on top of fine-grained data flow scheduling, each core also has 8 microthreads holding 8 independent tensor instructions. The scheduler chooses among these cores to execute in the core compute logic, guaranteeing that there are always useful work ready for compute logic to execute.
@@
Cores in Cererbas WSE are arranged in a 2-D mesh topology. Each core has a router that has a 32-bit bidirectional port to four adjacent cores in north, south, east, and west direction, as well as one 32-bit port to the Compute logic within the core. Data packets in Cerebras WSE are 32-bit long, 16-bit data and 16-bit control. Packets are communicated through static routing. Each router has 24 local static routes that can be configured, called colors. The static routing mechanism and 2D mesh topology enable high-speed data communication among cores [15].
@@
Cerebras WSE also uses weight streaming to enable training very large models. Model weights are stored in an external memory device of DRAM and flash memory called MemoryX. MemoryX sends the weights of each layer to Cerebras WSE at a time. After Cerebras WSE computes the output values using the streamed weights and the data in its cores, backpropagation occurs and the gradients of the layer are sent back to MemoryX to perform weight updates. Storing model weights externally enables extremely large model sizes as they do not consume on-chip memory [15].
@@ III-B Datasets
We utilized the PILE [17] dataset for training of GPT-3 models. The PILE is an 825.18 GB, containing 211, 043, 181 samples [26], English text dataset collected from 22 high-quality sub-datasets of different domains. By training LLMs on datasets of diverse domains, the model is able to gain greater general domain knowledge. We utilize the PILE dataset for pre-training of the GPT-3 models. The task completed by the model is to predict the next word based on all previous words in the sequence. Therefore, the label of each training sample is the input sequence tokens right shifted by one token. To simplify the training settings and reduce the time taken to complete the experiments, we use a 16MB subset, containing 10000 samples, of the PILE dataset for training of GPT-3 models.
@@ III-C Model Hyper-parameters
We performed our experiments on the BERT-base ( [^1] M) and BERT-large (340M) variants. For GPT-3, we evaluated the 256M, 590M, 2.7B, 6.7B, 13B, and 20B model sizes. Table II and Table III represent the hyper-parameters of each model variant. In these tables, $L$ represents the number of layers (transformer blocks). $H$ represents the hidden size. $A$ represents the number of self-attention heads. $MSL$ represents the maximum sequence length of the input.
@@ III-D Performance metrics
For training of BERT and GPT-3 models, we use throughput, samples per second, as the perofrmance metric. For inference of BERT models, we use end-to-end latency as the performance metric. Because BERT models perform classification tasks, we define the end-to-end latency of BERT models as the duration from when the input batch is given to the model to when the model output is produced.
@@ IV-B GPT-3 Training Performance Analysis
Training experiments of GPT-3 models were conducted on the GPT-3 model sizes indicated in Table III. Figure 5 shows the training throughput of each GPT-3 model size with varying batch sizes. For GPT-3 256M model, training throughput increases slightly as batch size increases. We believe that this is due to the model is able to take more advantage of the high computation resources on the Cerebras WSE by having large batch sizes because the model weights being transferred onto the chip consume less memory bandwidth and large batch sizes can enable more computation to occur simultaneously. For GPT-3 20B size, training throughput drops at large batch size. This could be potentially due to the fact that the memory bandwidth on the chip is fully utilized and data transfer dominates the training time. For other GPT-3 model sizes, the training throughout stays roughly the same across model sizes.
@@ IV-C BERT Inference Performance Analysis
Figure 6 shows the inference latency for BERT models over varying batch sizes. Our experiment results indicate that the BERT-base model latency does not vary by much for all batch sizes except batch size of 1. Likewise, BERT-large model latency does not change much after a certain batch size. However, recall that end-to-end latency is measured as the duration takes for the output of the entire batch to be produced. Therefore, this indicates that performing inference for larger batch size is still beneficial on the Cerebras WSE platform as the end-to-end latency does not vary by much with increasing batch size, therefore making the average inference latency per sample decrease with larger batch size.
@@ IV-D Roofline Models
Based on our experiment results, we completed roofline model analysis for BERT and GPT-3 training, shown in Figure 7 Figure 8. Our results show that both BERT-base and BERT-large training on all the batch sizes we investigated operate in the compute-bound region. Moreover, all GPT-3 training for batch size of 128 operate in the compute-bound region.
@@
Figure 8: GPT-3 roofline analysis. Axes are in log scale for better visualization. Only batch size of 128 is investigated
@@ V Conclusion and Future Work
In this paper, we examined the training throughput of BERT and GPT-3 training across commonly used batch sizes and model sizes, as well as inference latency for BERT. We also investigated the roofline model of BERT training, gaining insights in the scalability and potential of the Cerebras WSE for accelerating LLMs training and inference. Investigating Cerebras WSE’s performance for more models, especially computer vision models, will be a promising area for exploration because of its widespread applications.

### v2

@@ Abstract
In this paper, we evaluate the performance of LLMs on the Cerebras Wafer Scale Engine (WSE). Cerebras WSE is a high performance computing system with 2.6 trillion transistors, 850,000 cores and 40 GB on-chip memory. Cerebras WSE’s Sparse Linear Algebra Compute (SLAC) cores eliminate multiply-by-zero operations and its 40 GB of on-chip memory is uniformly distributed among SLAC cores, enabling fast local access to model parameters. Moreover, Cerebras software configures routing between cores at runtime, optimizing communication overhead among cores. As LLMs are becoming more ubiquitous used, new hardware architectures are required to accelerate LLM training and inference. We benchmark the effectiveness of this hardware architecture at accelerating LLM training and inference. Additionally, we analyze if Cerebras WSE can scale the memory-wall associated with traditionally memory-bound compute tasks using its 20 PB/s high bandwidth memory. Furthermore, we examine the performance scalability of Cerebras WSE through a roofline model. By plotting training throughput against computational intensity, we aim to assess their effectiveness at handling high compute-intensive LLM training and inference tasks.
@@ II-A Large Language Models (LLMs)
Large language models have gained growing interest over the past few years. Especially with the release of ChatGPT [19], users began recognizing the capabilities of LLMs and leverage them to assist with daily tasks that were previously time-consuming. Because of its remarkable language understanding abilities, more than one million users signed up to use ChatGPT within one week of its release [20]. Moreover, LLMs have created a profound influence in many industries and fields. For example, LLMs ability to generate and debug code can accelerate the software development process and enhance programmers’ productivity [21]. LLMs can also assist with the creative writing process by offering new ideas and providing feedback and grammatical corrections on input writings of users [20].
@@
$\bm{A}=\text{softmax}\left(\frac{\bm{Q}\bm{K}^{T}}{\sqrt{{d}_{k}}}\right)\bm{V}$ (4)
$\bm{W}^{Q}\in\mathbb{R}^{d\times{d}_{k}}$, $\bm{W}^{K}\in\mathbb{R}^{d\times{d}_{k}}$, $\bm{W}^{V}\in\mathbb{R}^{d\times{d}_{v}}$ are the weight matrices for query, key, and value, respectively. $\bm{Q}\in\mathbb{R}^{n\times{d}_{k}}$, $\bm{K}\in\mathbb{R}^{n\times{d}_{k}}$, $\bm{V}\in\mathbb{R}^{n\times{d}_{v}}$ are the query, key, and value matrices, respectively. $\bm{X}\in\mathbb{R}^{n\times d}$ is the input sentence embedding matrix, where $n$ is the input sequence length, $d$ is the embedding length of each word in the input sentence. The SA output $\bm{A}\in\mathbb{R}^{n\times d_{v}}$ will then be projected to the same shape as input $\bm{X}$ through a feed-forward network. This sequence of operations define the essential elements of a transformer block and is also called a single attention head. In transformer-based LLMs, each transformer block can contain multiple such attention heads. Each head captures different relationships among the tokens in the input sequence, baking within the model a richer context and understanding via the Multi-Headed Self-Attention (MHSA) mechanism. As the input and output dimensions of a transformer block are the same, multiple transformer blocks can be stacked, allowing earlier transformer blocks to provide more context for later blocks.
Transformer models are pre-trained over a large corpus of tokens via tasks such as next word prediction (GPT) [22] and masked language modeling (BERT) [23].
@@ II-B GPT
GPT is a transformer decoder-only architecture where each input only computes SA among previous input tokens. Namely, when computing output SA token ${\bm{a}_{i}\in\mathbb{R}^{d_{v}}}$, SA only computes among ${\bm{q}_{i}}\in\mathbb{R}^{d_{k}}$ with ${\bm{k}_{j}\in\mathbb{R}^{d_{k}}}$, ${\bm{v}_{j}\in\mathbb{R}^{d_{v}}}$ where $j\leq i$. This ensures that the model only attends to past and current tokens, not future unseen ones. At each time stamp, GPT computes a probability distribution over the entire vocabulary to predict the probability of next word, enabling it to generate the next word. This process is also called auto-regressive decoding or text generation. Auto-regressive text generation during inference terminates when an endoftext token is generated or a predefined maximum output sequence length is reached.
Figure 1: GPT transformer decoder architecture. Input embeddings are first added with positional embeddings and then fed into GPT transformer-decoder block, where MHSA is performed.
@@ II-C BERT
BERT[12] is a transformer encoder-only architecture where each input computes SA among both previous and future input tokens, called bidirectional SA. Namely, when computing output SA token ${\bm{a}_{i}\in\mathbb{R}^{d_{v}}}$, SA only computes among ${\bm{q}_{i}}\in\mathbb{R}^{d_{k}}$ with ${\bm{k}_{j}\in\mathbb{R}^{d_{k}}}$, ${\bm{v}_{j}\in\mathbb{R}^{d_{v}}}$ where $j\leq n$, where $n$ is the input sequence length. Because SA is computed over the entire input sequence, BERT models have a global comprehensive context over all tokens in the input sequence. Normally, a [class] token is added at the beginning of the input sequence. At the output layer, the final hidden state of the [class] token is used for the classification of the entire sentence. This ensures that the final classification of the sentence is not biased toward any specific word in the input sequence. BERT models are typically pre-trained on tasks like Masked Language Modeling (MLM) [12], where a random subset of the input sequence is masked and the model predicts the masked words. BERT models are also pre-trained via Next Sentence Prediction (NSP), where the model predicts whether the given two sentences should follow each other. After pre-training, BERT models are fine-tuned for task-specific applications, such as sentiment analysis of input sequence [24].
Figure 2: BERT transformer encoder architecture. [class] token embedding is appended to the front of the input sequence. The BERT architecture is similar to the GPT model architecture shown in Figure 1, except that MHSA occurs between the current token and tokens to the left and right. In BERT, the final output hidden state of [class] token is used to predict the probability distribution for classification of the input sequence.
@@ II-D Cerebras WSE
All training and inference experiments are conducted on the Cerebras Wafer Scale Engine (WSE) and Cerebras CS-2 system. Cerebras WSE is a powerful AI accelerator built to train LLMs [25]. As of May 2023, Cerebras WSE is the largest chip ever built, at 46,000 mm2 and containing 2.6 trillion transistors and 850,000 cores. Cerebras WSE’s large die size enables greater compute power and enables more computations to occur on-chip [16]. This reduces off-chip communications and enables computation resources to be fully utilized. Cerebras chip is built with Taiwan Semiconductor Manufacturing Company (TSMC) 7 nm process and runs at 1.1 GHz frequency [15]. The WSE’s physical design contributes 50% silicon area to static random access memory (SRAM) and 50% to computation logic in each core. Each core features a local 48 KB SRAM that enables the core to have fast access to local memory. In each core, memory is organized into eight 6 KB banks, each 32-bit wide. On top of the local 48 KB memory, each core also has a 256-byte, 64-bit wide, software-configured cache, enabling even faster data access for most frequently used data [15]. This uniformed distributed on-chip memory architecture enables low memory access latency for each core and very high aggregated memory throughput.
@@
Cerebras WSE’s fine-grained data flow scheduling, enables cores to only perform computations on non-zero data, saving dynamic power of cores [15]. The combination of this and the high memory bandwidth enables efficient computations on data with unstructured sparsity [26], which is well suited for neural network computations as model weights and input data can possess arbitrary levels of sparsity. Moreover, on top of fine-grained data flow scheduling, each core also has 8 micro-threads holding 8 independent tensor instructions. The scheduler chooses among these cores to execute in the core compute logic, guaranteeing that there is always useful work ready for the compute logic to execute.
@@
Cores in Cerebas WSE are arranged in a 2-D mesh topology. Each core has a router that has a 32-bit bidirectional port to four adjacent cores in north, south, east, and west direction, as well as one 32-bit port to the compute logic within the core. Data packets in Cerebras WSE are 32-bit long, 16-bit data and 16-bit control. Packets are communicated through static routing. Each router has 24 local static routes that can be configured, called colors. The static routing mechanism and 2D mesh topology enable high-speed data communication among cores [15].
@@
Cerebras WSE also uses weight streaming to enable training very large models. Model weights are stored in an external memory device DRAM and flash memory called MemoryX. MemoryX sends the weights of each layer to Cerebras WSE at compute time. After Cerebras WSE computes the output values using the streamed weights and the data in its cores, backpropagation occurs and the gradients of the layer are sent back to MemoryX to perform weight updates. Storing model weights externally enables extremely large model sizes as they do not consume on-chip memory [15].
@@ III-B Datasets
We utilized the PILE [17] dataset for training of GPT-3 models. The PILE is an 825 GB dataset, containing 211,043,181 samples [27] of English text dataset collected from 22 high-quality sub-datasets of different domains. By training LLMs on datasets of diverse domains, the model is able to gain greater general domain knowledge. We utilize the PILE dataset for pre-training of the GPT-3 models. The task completed by the model is to predict the next word based on all previous words in the sequence. Therefore, the label of each training sample is the input sequence tokens right shifted by one token. To simplify the training settings and reduce the time taken to complete the experiments, we use a 16 MB subset, containing 10,000 samples, of the PILE dataset for training of the GPT-3 models.
@@ III-C Model Hyper-parameters
We performed our experiments on the BERT-base ( [^1] M) and BERT-large (340M) variants. For GPT-3, we evaluated the 256M, 590M, 2.7B, 6.7B, 13B, and 20B model sizes. Table II and Table III represent the hyper-parameters of each model variant. In these tables, $L$ represents the number of layers (transformer blocks), $D$ represents the hidden size, $H$ represents the number of self-attention heads and $MSL$ represents the maximum sequence length of the input.
@@ III-D Performance metrics
For training of BERT and GPT-3 models, we use throughput, samples per second, as the performance metric. For inference of BERT models, we use end-to-end latency as the performance metric. Because BERT models perform classification tasks, we define the end-to-end latency of BERT models as the duration from when the input batch is given to the model to when the model output is produced.
@@ IV-B GPT-3 Training Performance Analysis
Training experiments of GPT-3 models were conducted on the GPT-3 model sizes indicated in Table III. Figure 5 shows the training throughput of each GPT-3 model size with varying batch sizes. For GPT-3 256M model, the training throughput increases as the batch size increases. Larger batch sizes can more effectively utilize the available high bandwidth memory. This facilitates more computations within each core, exploiting the large compute resources available in the WSE. For GPT-3 20B size, training throughput drops at large batch sizes. This is due to the fact that the memory bandwidth on the chip is fully utilized and data transfer dominates the compute time during training. For other GPT-3 model sizes, the training throughput stays roughly the same across model sizes. This highlights the WSE’s unique capacity to scale training for large models and batch sizes without a drop in throughput.
@@ IV-C BERT Inference Performance Analysis
Figure 6 shows the inference latency for BERT models over varying batch sizes. Our experiment results indicate that the BERT-base model latency does not vary by much for all batch sizes except batch size of 1. Likewise, BERT-large model latency does not change much after a certain batch size. However, recall that end-to-end latency is measured as the duration it takes for the output of the entire batch to be produced. Therefore, this indicates that performing inference for larger batch sizes is beneficial on the Cerebras WSE platform as the end-to-end latency does not vary by much with increasing batch size. This reduces the average inference latency per sample for larger batch sizes, as is expected for high-compute high-bandwidth systems like Cerebras.
@@ IV-D Roofline Models
Based on our experimental results, we completed roofline model analysis for BERT and GPT-3 training, shown in Figure 7 and Figure 8. Our results show that both BERT-base and BERT-large training on all the batch sizes we investigated operate in the compute-bound region. Moreover, all GPT-3 training for batch size of 128 operate in the compute-bound region. This highlights that training large models on the Cerebras WSE is scalable, with Cerebras WSE scaling the memory wall for LLM training through its unique architecture.
@@
Figure 8: GPT-3 roofline analysis. Axes are in log scale. Only batch size of 128 is investigated.
@@ V Conclusion and Future Work
In this paper, we examined the training throughput of BERT and GPT-3 training across commonly used batch sizes and model sizes, as well as inference latency for BERT. We also investigated the roofline model of BERT and GPT-3 training, gaining insights in the scalability and potential of the Cerebras WSE for accelerating LLM training and inference. Investigating Cerebras WSE’s performance for more models, especially computer vision models, will be a promising area for exploration because of its widespread applications.